7. On a serve, a tennis ball may leave the tennis racket of a top player at a speed of 65.0 m/s. If the ball has a mass of 0.0600 kg and it is in contact with the racket for 0.0300 seconds, what is the average force applied to the ball? Would that be enough force to lift an average size person (75 kg)? Assume that the tennis ball is initially not moving
Answer #1

By impulse-momentum theorem,

Impulse = change in momentum

=> F*t = mv

=> F = mv/t

=> F = 0.0600*65/0.0300

= 130 N (answer)

Weight of the person = mg = 75*9.8 = 735 N

F is less than the weight so it is not enough to lift that person.